OpenStudy (anonymous):

What is the common ratio of the following geometric sequence? 4/15, 2/5, 3/5, 9/10,27/20 A. 2/3 B. 3/2 C. 1/5 D. 4/15

OpenStudy (anonymous):

@Godlovesme or @dtan5457 can someone help?

OpenStudy (campbell_st):

same as before 4/15 x r = 2/5 to check is it the same as 2/5 x r = 3/5

OpenStudy (godlovesme):

^^^^^ @Tallan